OpenStudy (anonymous):

How do you work out 33 1/3 of 27

OpenStudy (anonymous):

33 1/3 = 100/3 ; 100/3 X 27 =900

OpenStudy (anonymous):

( (33*3) +1 ) / 3 = ( 100 * 27 ) / 3 = 2700 / 3 = 900